# Free-style and Fast 3D Portrait Synthesis

### v1

@@ Abstract
Efficiently generating a free-style 3D portrait with high quality and consistency is a promising yet challenging task. The portrait styles generated by most existing methods are usually restricted by their 3D generators, which are learned in specific facial datasets, such as FFHQ. To get a free-style 3D portrait, one can build a large-scale multi-style database to retrain the 3D generator, or use a off-the-shelf tool to do the style translation. However, the former is time-consuming due to data collection and training process, the latter may destroy the multi-view consistency. To tackle this problem, we propose a fast 3D portrait synthesis framework in this paper, which enable one to use text prompts to specify styles. Specifically, for a given portrait style, we first leverage two generative priors, a 3D-aware GAN generator (EG3D) and a text-guided image editor (Ip2p), to quickly construct a few-shot training set, where the inference process of Ip2p is optimized to make editing more stable. Then we replace original triplane generator of EG3D with a Image-to-Triplane (I2T) module for two purposes: 1) getting rid of the style constraints of pre-trained EG3D by fine-tuning I2T on the few-shot dataset; 2) improving training efficiency by fixing all parts of EG3D except I2T. Furthermore, we construct a multi-style and multi-identity 3D portrait database to demonstrate the scalability and generalization of our method. Experimental results show that our method is capable of synthesizing high-quality 3D portraits with specified styles in a few minutes, outperforming the state-of-the-art. Project page is available at https://tianxiangma.github.io/FF3D.
@@ 1 Introduction
Recently, 3D portrait synthesis has attracted more and more attention, especially with the emergence of Neural Radiance Field (NeRF) [7]. As the representatives among them, 3D-aware GAN methods [3, 4, 8] combine NeRF with StyleGANs [2] to ensure 3D consistency synthesis. By mapping an image to the 3D GAN latent space, 3D GAN inversion approaches [9, 10, 11] can generate or edit a specific 3D portrait. However, both of them fail to create a free-style 3D portrait, e.g., a style defined by user’s text prompt, since their generators are usually trained on a dataset that follows a particular style distribution, such as the realism style in FFHQ [1], which raises a question: how to generate a free-style 3D portrait at a low cost? One may collect a large number of portrait images with different styles to retrain their models, but the data preparing and training process are usually time-consuming. Another potential solution is that synthesizing a style-specific 3D portrait first, then transferring it to any style with a off-the-shelf style transfer tool. Unfortunately, the 3D consistency will be difficult to be maintained.
Figure 1: The 3D portrait synthesis results based on single-style few-shot dataset. The first row shows man with different styles, the second row shows various characters, and the last two rows show diverse identities with different styles from varied viewpoints.
@@
To this end, we propose an efficient pipeline to achieve free-style 3D portrait synthesis in this paper. At first, we leverage the knowledge of two pre-trained generative priors, EG3D [4] and Instruct-pix2pix (Ip2p)[12] to construct a few-shot portrait dataset with a given style, avoiding dirty data collection and cleaning. The former generates a multi-view 3D portrait, and the latter performs text-guided style editing in each viewpoint. We empirically find the editing results of Ip2p vary significantly along viewpoints for some specified style, resulting in the issue of multi-view misalignment. To alleviate this problem, an optimization strategy is introduced into the inference stage of Ip2p. Secondly, we replace the original triplane generator with a trainable I2T network, and freeze other parts of EG3D to avoid training the total EG3D. We pre-train the I2T network to create the mapping from portrait image to triplane feature space, which helps I2T network fine-tune the few-shot set in a few minutes. Our high-quality stylized 3D portrait synthesis results are shown in Fig. 1. At last, our approach is capable of synthesizing multiple styles as well as multiple identities with the help of the multi-style and multi-identity portrait dataset that we construct previously, demonstrating its potential scalability and generalization.
@@ 2.2 Text-guided image editing
There are numerous image editing methods, and the performance of text-guided image editing methods [29, 30, 31, 12] has been qualitatively improved thanks to the advancement of pre-trained image generation large models [32, 33, 34] based on the Diffusion model. Ip2p [12] is a SOTA text-guided image editing method, which uses two generative priors, GPT-3 [35] and Stable Diffusion [32], to synthesize a large number of paired images and then train a conditional diffusion model on them. This model allows the users to provide a relatively free text instruction to edit a given image, including stylistic transfer. Therefore, Ip2p is well suitable as a text-guided image editing prior for this paper to perform text-guided style transfer on portraits from different viewpoints. However, the model also has problems, such as poor generation with some simple text prompts and generating portraits with large stylistic variations for different views of the same portrait. We propose some improvements to solve these problems in this paper.
Figure 2: The framework of our method. We introduce two generative priors EG3D and Ip2p to quickly build a few-shot stylized portrait dataset $\mathcal{D}_{s}$. Ip2p is optimized to produce more stable and consistent text-guided stylization results. With the constructed few-shot dataset, we quickly fine-tune a pre-trained I2T network to achieve stylized 3D portrait synthesis.
@@ 3 Methodology
In this section, we detail our free-style and fast 3D portrait synthesis framework, as shown in Fig. 2. We first briefly introduce two generative priors, EG3D and Ip2p, and combine them to build a few-shot dataset with a given style. To describe styles more freely, we optimize Ip2p to make it more stable for stylizing portraits from different perspectives (Sec. 3.1). We then use the few-shot dataset to fine-tune our proposed I2T network, which is equipped with EG3D prior to achieve fast stylized 3D portrait synthesis (Sec. 3.2).
3.1 Few-shot dataset construction
We denote the input style as $\mathbf{t}$, which is often defined by text prompt from user.
@@ 3.1 Few-shot dataset construction
Text-guided image editing prior. Next, we employ Ip2p [12] to perform the style editing on the portraits generated above. Ip2p implements conditional image editing based on Stable Diffusion [32], which can be denoted as $\mathit{T}(\phi,\mathit{I},\mathbf{n},\mathbf{c})$, where $\phi$ is the model parameters, $\mathit{I}$ is the input portrait, $\mathbf{n}$ is the Gaussian noise used in the denoising process, and $\mathbf{c}$ is the text prompt that guides the editing direction. Considering different input noise will generate different results, we fix $\mathbf{n}$ to keep the identity unchanged. Then we rewrite $\mathbf{t}$ as $\mathbf{c}$, let $\mathit{I}$ be the portraits that are produced by EG3D, and use $\mathit{T}$ to generate stylized portraits.
@@
where $\mathit{I_{s}}$ is one stylized portrait. We construct a few-shot stylized portrait dataset $\mathcal{D}_{s}$ using different $\mathbf{v}$ from $(P,Y)$, so $\mathcal{D}_{s}$ contains $i^{2}$ stylized portraits. The construction pipeline is summarized in Algorithm 1.
3.2 Image-to-Triplane network
Considering the ability in generating high quality and 3D consistency image, we also utilize EG3D to synthesize 3D portrait. With the few-shot dataset, an intuitive solution is using it to fine-tune a pre-trained EG3D. However, we find that training the entire EG3D is not efficient, and the rich facial prior hidden in the pre-trained EG3D can also be damaged. In order to fully enjoy the priors, one may invert $\mathcal{D}_{s}$ to the $\mathcal{W}$ latent space. Unfortunately, it will limit the portrait style within the scope of training set that EG3D is pre-trained, since the triplane feature generator is not changed. As a result, we introduce I2T module to learn the mapping from $\mathcal{D}_{s}$ to triplane feature space. We replace the original triplane generator of EG3D with our I2T module and keep other parameters unchanged as a EG3D renderer, reaching a trade-off between training efficiency and prior utilization.
Similar to the triplane generation network of the original EG3D, I2T consists of multiple StyleGAN modulation layers (SMLs) and a style encoder $E$, as shown at the bottom left corner of Fig. 2. Due to the small size of $\mathcal{D}_{s}$, we expand the latent code $\mathbf{w}_{s}$ along spatial dimension (from $c$ to $c\times k\times k$) to enrich it with style and structural information. Then we adopt $E$ to learn $\mathbf{w}_{s}$ first, and feed it into SMLs to generate stylized triplane. For a particular feature layer $F^{i}\in\mathbb{R}^{C_{i}\times H_{i}\times W_{i}}$ in SMLs, we have the following formula:
@@ 3.2 Image-to-Triplane network
where $\mu_{h,w}^{i}$ and $\sigma_{h,w}^{i}$ represent the calculated mean and standard deviation across channel dimension, respectively. $\gamma_{h,w}^{i}$ and $\beta_{h,w}^{i}$ are learnable weight networks. It is noted that, to ensure stable feature learning, the input portrait of $E$ is always of front view, which contains richest style information compared with other viewpoint.
Although I2T network can be trained on $\mathcal{D}_{s}$, it still suffers from two challenges: First, $\mathcal{D}_{s}$ is a few-shot dataset, it has no more than 100 images for a style in practice. Second, the portraits in $\mathcal{D}_{s}$ have more or less differences in style, resulting in 3D inconsistency. We alleviate this problem by pre-training the I2T network on the portraits generated by EG3D prior. In particular, in each iteration of the pre-training, we randomly generate a portrait from EG3D, and record its triplane representation as $\mathbf{p}\in\mathbb{R}^{256\times 256\times 96}$, its front view as $\mathit{I_{f}}\in\mathbb{R}^{512\times 512\times 3}$. Then the I2T network is pre-trained using the following loss function:
$\mathcal{L}_{\textrm{I2T}}=\mathbb{E}_{\mathit{I_{f}},\mathbf{p}}[\left\|\mathit{H}(\mathit{I_{f}})-\mathbf{p}\right\|_{1}],$ (4)
where $\mathit{H}$ represents the I2T network. After some iterations, we can learn the mapping between the input portrait and its triplane representation.
@@
Algorithm 2 The I2T network training
1:Input: $\mathcal{D}_{s}$, Pre-trained I2T
2:Init I2T net using the pre-trained I2T
@@
4:  Select $\mathit{I_{v}},\mathit{I_{f}}$ from $\mathcal{D}_{s}$
5:  Fine-tuning I2T net in $\mathcal{D}_{s}$ using loss:
@@
Training. After being pre-trained, the I2T network can be quickly fine-tuned on the stylized multi-view portrait dataset $\mathcal{D}_{s}$, as shown in the lower part of Fig. 2. For an arbitrary view portrait $\mathit{I_{v}}$ and front view portrait $\mathit{I_{f}}$ in $\mathcal{D}_{s}$, we use the following image reconstruction loss:
$\mathcal{L}_{\textrm{rec}}=\mathbb{E}_{\mathit{I_{v}},\mathit{I_{f}}\in\mathcal{D}_{s}}[\left\|\mathit{G^{*}}(\mathit{I_{f}},\mathbf{v})-\mathit{I_{v}}\right\|_{1}+lpips(\mathit{G^{*}}(\mathit{I_{f}},\mathbf{v}),\mathit{I_{v}})],$ (5)
where the former is the $L_{1}$ reconstruction loss, $\mathit{G^{*}}$ represents the new EG3D model with I2T network. $lpips(\cdot,\cdot)$ is the Learned Perceptual Image Patch Similarity [38] loss, which calculates the distance of the latent features extracted from the VGG network. In addition to the reconstruction loss, we add the density regularization, which encourages smoothness of the density field rendered by triplane and prevents sharp or hollow portrait shapes during fine-tuning. The density regularization loss is shown as follows,
@@
where $\lambda_{\textrm{rec}}$ and $\lambda_{\textrm{dr}}$ are loss weights.
Thanks to the pre-trained I2T, we can quickly extend the domain of I2T to stylized portraits $\mathcal{D}_{s}$ with only a few training iterations. The total training process is listed in Algorithm 2.
@@ 4.1 Implementation detail
Our method is trained on a single NVIDIA A100, using the Adam optimizer with learning rate of $0.002$ and $\beta_{1}=0,\beta_{2}=0.99$. In Ip2p inference, we set the number of time step $\mathbf{s}$ in DDIM [39] to 20. The degree of noise addition $\tau$ is 0.9, and the image guide and text guide weight parameters of Ip2p are set to 1.5 and 20.0. The spatial dimension $k$ of $\mathbf{w}_{s}$ is 32. When fine-tuning the I2T network, the loss weights are set as $\lambda_{\textrm{rec}}=10.0$ and $\lambda_{\textrm{dr}}=0.2$. The number of samples $\mathbf{i}$ is 10. Other parameters, such as camera focal length, use the EG3D default settings. The decorative prompt $\mathbf{t_{d}}$ is "realistic, detail, 8k, photorealistic", then the positive prompt input for Ip2p is "turn the head into $\mathbf{t}$, $\mathbf{t_{d}}$". The negative prompt $\mathbf{t_{n}}$ is "unclear facial features, non-face objects, ugly, bad". Note that we fix $\mathbf{t_{d}}$ and $\mathbf{t_{n}}$, and only change $\mathbf{t}$ in all following experiments. Fine-tuning $\mathbf{t_{d}}$ or $\mathbf{t_{n}}$ will polish the generated results, but it’s not the scope of this paper.
Figure 3: Multi-style 3D portrait synthesis using single model.
Figure 4: Multi-style and multi-identity 3D portrait synthesis using single model.
4.2 Multi-style and multi-identity portrait dataset
We also construct a multi-style and multi-identity portrait dataset to demonstrate our scalability and generalization. ChatGPT [40] is adopted to generate text prompts for describing portrait styles.
Multi-style and single-identity portrait dataset (MSSI). We use ChatGPT to generate 100 categories of portrait style description prompts, including art style, game character style, movie character style and basic attribute editing. For each text prompt, we use Algorithm 1 to synthesize 100 stylized portraits from different viewpoints $(10\ pitch\times 10\ yaw)$. So the multi-view and multi-style dataset consists of 10,000 portraits in total.
Multi-style and multi-identity portrait dataset (MSMI). The previous dataset only contains one identity for all portraits, i.e., each portrait is obtained from the same $\mathbf{w}$ vector in Algorithm 1. In order to explore multi-style and multi-identity 3D portrait synthesis, we reuse the above 100 style prompts, but sample 100 different $\mathbf{w}$ vectors in each style (one $\mathbf{w}$ for one viewpoint). The multi-style and multi-identity dataset also has 10,000 portraits.
Fine-tuning our model on MSSI and MSMI will enable multi-styled and multi-identity portrait synthesis with a single model.
4.3 Free-style 3D portrait synthesis
@@
In this section, we show the 3D portrait synthesis results of our approach. As shown in Fig. 1, our method can generate diverse style 3D portrait, and the synthesized portraits are high quality and 3D consistent, proving our ability of free-style generation. In addition, we conduct multi-style and multi-identity 3D portrait synthesis experiments on MSSI and MSMI. We select 20 styles and display 5 viewpoints for one portrait in Fig. 3, and show 8 identities for one style in Fig. 4, it is clear our method is capable of synthesizing high fidelity multi-style and multi-identity 3D portrait. Moreover, MSSI and MSMI can be trained with a single model, and they can be expanded to larger style or identity size, demonstrating the effectiveness and generalization of our method. Last but not least, Fig. 5 shows the smooth style mixing results of our method. We interpolate the encoded features of two different style in the $\mathbf{w}_{s}$ latent space of the model trained on MSSI, and use the new triplane representation to synthesize 3D portraits. The results reveal the controllability and potential style editing ability of our model.

### v3

# Freestyle 3D-Aware Portrait Synthesis Based on Compositional Generative Priors

@@ Abstract
Efficiently generating a freestyle 3D portrait with high quality and 3D-consistency is a promising yet challenging task. The portrait styles generated by most existing methods are usually restricted by their 3D generators, which are learned in specific facial datasets, such as FFHQ. To get the diverse 3D portraits, one can build a large-scale multi-style database to retrain a 3D-aware generator, or use a off-the-shelf tool to do the style translation. However, the former is time-consuming due to data collection and training process, the latter may destroy the multi-view consistency. To tackle this problem, we propose a novel text-driven 3D-aware portrait synthesis framework that can generate out-of-distribution portrait styles. Specifically, for a given portrait style prompt, we first composite two generative priors, a 3D-aware GAN generator and a text-guided image editor, to quickly construct a few-shot stylized portrait set. Then we map the special style domain of this set to our proposed 3D latent feature generator and obtain a 3D representation containing the given style information. Finally we use a pre-trained 3D renderer to generate view-consistent stylized portraits from the 3D representation. Extensive experimental results show that our method is capable of synthesizing high-quality 3D portraits with specified styles in a few minutes, outperforming the state-of-the-art.
Figure 1: Our freestyle 3D-aware portrait synthesis results. The first row shows man with different styles, the second row shows various characters, and the last two rows show diverse identities with different styles from varied viewpoints. Zoom in for better viewing.
$*{}^{*}$ Corresponding author.
@@ 1 Introduction
Recently, 3D portrait synthesis has attracted more and more attention, especially with the emergence of Neural Radiance Field (NeRF) [29]. As the representatives among them, 3D-aware GAN methods [13, 7, 33] combine NeRF with StyleGANs [20] to ensure 3D consistency synthesis. By mapping an image to the 3D GAN latent space, 3D GAN inversion approaches [23, 24, 43] can generate or edit a specific 3D portrait. However, both of them fail to create a freestyle 3D portrait, e.g., a style defined by user’s text prompt, since their generators are usually trained on a dataset that follows a particular style distribution, such as the realism style in FFHQ [19], which raises a question: how to generate a freestyle 3D portrait at a low cost? One may collect a large number of portrait images with different styles to retrain their 3D generative models, but the data preparing and training process are usually time-consuming. Another potential solution is that synthesizing a dataset-based 3D portrait first, then transferring each portrait view to other styles with a off-the-shelf 2D style transfer tool. Unfortunately, the 3D consistency will be difficult to be maintained.
@@
To this end, we propose an efficient framework to achieve freestyle 3D-aware portrait synthesis in this paper. At first, we composite the knowledge of two pre-trained generative priors, EG3D [7] and Instruct-pix2pix (Ip2p) [3] we used in this paper, to construct a few-shot stylized portrait dataset with a given style prompt, avoiding dirty data collection and cleaning. The former generates a multi-view 3D portrait, and the latter performs text-guided style editing in each viewpoint. We empirically find the editing results of Ip2p vary significantly along viewpoints for some given style, resulting in the issue of multi-view misalignment. To alleviate this problem, some optimization strategies are introduced into the inference stage of Ip2p. Secondly, we propose a 3D latent feature generator model, which can map the style information from stylized portrait dataset to 3D latent representation. This generator is first pre-trained with large amount of generative data from EG3D, so that it can learn 3D latent representation well from multi-view portraits. Our 3D latent feature generator enables fast fine-tuning with few-shot stylized portrait dataset, and generates out-of-distribution stylized 3D representation. Finally, we utilize the pre-trained EG3D neural renderer and super-res generator as our 3D Renderer, and we find that this model does not need to be fine-tuned to generate 3D-consistent portrait with a specified style when we have learned a well-stylized 3D implicit representation. Our high-quality stylized 3D-aware portrait synthesis results are shown in Fig. 1, each style is specified by a text prompt, and each 3D portrait model can be fine-tuned in 3 minutes. Our main contributions can be summarized as follows:
  • We propose a 3D-aware portrait synthesis framework based on compositional generative priors that can be text-driven to generate freestyle 3D avatar.
  • We further propose a 3D latent feature generator, which allows for quick fine-tuning to map out-of-distribution style to 3D representation.
  • Compared with stylized 3D portrait synthesis baselines, our approach has clear advantages in performance and efficiency.
Figure 2: The framework of our method. We composite two generative priors EG3D and Ip2p, a 3D-aware GAN generator and a text-guided image editor, to quickly build a few-shot stylized portrait dataset $\mathcal{D}_{s}$, whose style is specified by the text prompt. Ip2p is optimized to produce more stable and consistent text-guided stylization results. With the constructed few-shot dataset, we quickly fine-tune a 3D latent feature generator to map the style information from the dataset into 3D representation, and then utilize a pre-trained 3D renderer to generate the stylized 3D-consistent portraits.
@@ 2.2 Text-guided image editing
There are numerous image editing methods, and the performance of text-guided image editing methods [2, 15, 22, 3, 14, 25] has been qualitatively improved thanks to the advancement of pre-trained image generation large models [38, 36, 37] based on the Diffusion model. Ip2p [3] is a SOTA text-guided image editing method, which uses two generative priors, GPT-3 [4] and Stable Diffusion [38], to synthesize a large number of paired images and then train a conditional diffusion model on them. This model allows the users to provide a relatively free text instruction to edit a given image, including stylistic transfer. Therefore, Ip2p is well suitable as a text-guided image editing prior for this paper to perform text-guided style transfer on portraits from different viewpoints. However, the model also has problems, such as poor generation with some simple text prompts and generating portraits with large stylistic variations for different views of the same portrait. We introduce some optimization strategies to alleviate these problems in this paper.
@@ 3 Methodology
In this section, we detail our freestyle 3D-aware portrait synthesis framework, as shown in Fig. 2. We first briefly introduce two generative priors, EG3D and Ip2p, and composite them to build a few-shot dataset with a given style. To describe styles more freely, we optimize the inference of Ip2p to make it more stable for stylizing portraits from different perspectives (Sec. 3.1). We then use the few-shot dataset to quickly fine-tune our proposed 3D latent feature generator, which is equipped with a pre-trained 3D renderer to generate 3D-consistent stylized portraits (Sec. 3.2).
3.1 Few-shot stylized portrait dataset construction
@@
Text-guided image editing prior. Next, we employ Instruct-pix2pix (Ip2p) [3] to perform the style editing on the portraits generated above. Ip2p implements conditional image editing based on Stable Diffusion [38], which can be denoted as $\mathit{T}(\phi,\mathit{I},\mathbf{n},\mathbf{t})$, where $\phi$ is the model parameters, $\mathit{I}$ is the input portrait, $\mathbf{n}$ is the Gaussian noise used in the denoising process, and $\mathbf{t}$ is the text prompt that guides the editing direction. Considering different input noise will generate different results, we fix $\mathbf{n}$ to keep the identity unchanged. Then we let $\mathit{I}$ be the multi-view portraits that are produced by EG3D, and use $\mathit{T}$ to generate stylized portraits.
@@
where $\mathit{I_{s}}$ is one stylized portrait. We construct a few-shot stylized portrait dataset $\mathcal{D}_{s}$ using different view $\mathbf{v}$ from $(P,Y)$, so $\mathcal{D}_{s}$ contains $i^{2}$ stylized portraits. The construction pipeline is summarized in Algorithm 1.
3.2 3D Latent Feature Generator
Considering the ability in generating high quality and 3D consistency image, we also utilize EG3D baseline to synthesize stylized 3D portrait. In the experiment, we find that if we have learned a well-stylized 3D implicit representation, the EG3D neural renderer and super-res generator do not need to be fine-tuned to generate 3D-consistent portrait with an out-of-distribution style. Therefore, we utilize the pre-trained EG3D neural renderer and super-res generator as our 3D renderer, and just need to learn a 3D latent feature generator to map the style information from $\mathcal{D}_{s}$ to the 3D implicit representation.
There are two other solutions that need to be discussed: 1) Directly fine-tuning the whole EG3D on $\mathcal{D}_{s}$, which is very inefficient compared to our method and generates somewhat blurred images. 2) Inverting the images from $\mathcal{D}_{s}$ to the $\mathcal{W}$ latent space of EG3D with GAN inversion methods, which will limit the portrait style within the scope of training set that EG3D is pre-trained, since the triplane feature generator is not changed.
We design the 3D latent feature generator as shown in the Fig. 2. An Encoder is used to learn the modulation latent feature $\mathbf{w}_{s}$ from different viewpoints, we expand the latent code $\mathbf{w}_{s}$ along spatial dimension (from $c$ to $c\times k\times k$) to enrich it with style and structural information. Then we propose the SMLs consisting of multiple Stylized modulation layers, which can learn to generate stylized triplane implicit representation. For a particular modulation layer in SMLs, we have the following formula:
@@
where $\mu_{h,w}^{i}$ and $\sigma_{h,w}^{i}$ represent the calculated mean and standard deviation across channel dimension, respectively. $\gamma_{h,w}^{i}$ and $\beta_{h,w}^{i}$ are learnable weight networks. The last layer of SMLs outputs the stylized triplane representation.
Although the 3D latent feature generator can be trained on $\mathcal{D}_{s}$, it still suffers from two challenges: First, $\mathcal{D}_{s}$ is a few-shot dataset, it has no more than 100 images for a style in practice. Second, the portraits in $\mathcal{D}_{s}$ have more or less differences in details, resulting in 3D inconsistency. We solve this problem by pre-training the 3D latent feature generator on the large number of portraits generated by EG3D prior, which can help to build accurate 3D implicit representation from multi-view portraits. In particular, in each iteration of the pre-training, we randomly sample to generate an arbitrary viewpoint portrait $\mathit{I_{v}}\in\mathbb{R}^{512\times 512\times 3}$ from EG3D, and record its triplane representation as $\mathbf{p}\in\mathbb{R}^{256\times 256\times 96}$, Then the pre-training process is constrained by the following loss function:
$\mathcal{L}_{\textrm{pre}}=\mathbb{E}_{\mathit{I_{v}},\mathbf{p}}[\left\|\mathit{H}(\mathit{I_{v}})-\mathbf{p}\right\|_{1}],$ (4)
where $\mathit{H}$ represents our 3D latent feature generator.
@@
Algorithm 2 The 3D latent feature generator fine-tuning
1:Input: $\mathcal{D}_{s}$, Pre-trained 3D latent feature generator
2:Init 3D generator net using the pre-trained model
@@
4:  Randomly select $\mathit{I^{1}_{v}},\mathit{I^{2}_{v}}$ from $\mathcal{D}_{s}$
5:  Fine-tuning 3D latent feature generator using loss:
@@
Figure 3: Multi-style and Multi-identity 3D portrait synthesis results. Zoom in for better viewing.
Figure 4: The outputs of different stages in our method.
Figure 5: Visualization of 3D geometry from different stylized generation results.
Training. After being pre-trained, the 3D latent feature generator can be quickly fine-tuned on the few-shot stylized portrait dataset $\mathcal{D}_{s}$. We use a multi-view cross-construction strategy for learning a more robust 3D implicit representation. For two arbitrary viewpoint portraits $\mathit{I^{1}_{v}}$ and $\mathit{I^{2}_{v}}$ in $\mathcal{D}_{s}$, we use the following reconstruction loss:
$\mathcal{L}_{\textrm{rec}}=\mathbb{E}_{\mathit{I^{1}_{v}},\mathit{I^{2}_{v}}\in\mathcal{D}_{s}}$ $[\left\|\mathit{G^{*}}(\mathit{I^{1}_{v}},\mathbf{v_{2}})-\mathit{I^{2}_{v}}\right\|_{1}$ (5)
$+lpips(\mathit{G^{*}}(\mathit{I^{1}_{v}},\mathbf{v_{2}}),\mathit{I^{2}_{v}})],$
where $\left\|\cdot\right\|_{1}$ is the $L_{1}$ reconstruction loss, $\mathit{G^{*}}$ represents the whole model with 3D latent feature generator and 3D renderer, and the parameters are input image and the rendering viewpoint, respectively. $lpips(\cdot,\cdot)$ is the Learned Perceptual Image Patch Similarity [44] loss, which calculates the distance of the latent features extracted from the VGG network. In addition to the reconstruction loss, we add the density regularization, which encourages smoothness of the density field rendered by triplane and prevents sharp or hollow portrait shapes during fine-tuning. The density regularization loss is shown as follows,
@@
where $\lambda_{\textrm{rec}}$ and $\lambda_{\textrm{dr}}$ are loss weights. The fine-tuning process is listed in Algorithm 2.
@@ 4.1 Implementation detail
Our method is implemented in PyTorch using an NVIDIA A100. We use Adam optimizer with learning rate of $0.002$ and $\beta_{1}=0,\beta_{2}=0.99$. The number of samples $i$ in few-shot dataset construction is 10. Other parameters, such as camera focal length, use the EG3D default settings. In Ip2p inference, we set the number of time step $\mathbf{s}$ in DDIM [40] to 20. The degree of noise addition $\tau$ is 0.9, and the image guide and text guide weight parameters of Ip2p are set to 1.5 and 20.0. The decorative prompt $\mathbf{t_{d}}$ is “realistic, detail, 8k, photorealistic”, then the positive prompt input for Ip2p is “turn the head into $\mathbf{t}$, $\mathbf{t_{d}}$”. The negative prompt $\mathbf{t_{n}}$ is “unclear facial features, non-face objects, ugly, bad”. Note that we fix $\mathbf{t_{d}}$ and $\mathbf{t_{n}}$, and only change $\mathbf{t}$ in all following experiments. Fine-tuning $\mathbf{t_{d}}$ or $\mathbf{t_{n}}$ will polish the generated results, but it’s not the scope of this paper. For each text prompt $\mathbf{t}$ we randomly sample a $\mathbf{w}$ and construct the few-shot dataset according to Algorithm 1. In 3D latent feature generator, The spatial dimension $k$ of $\mathbf{w}_{s}$ is 32. SMLs consists of 7 modulation layers. The 3D generator is pre-trained on EG3D randomly sampled data for 100k iterations. When fine-tuning the model on the few-shot dataset $\mathcal{D}_{s}$, the loss weights are set as $\lambda_{\textrm{rec}}=10.0$ and $\lambda_{\textrm{dr}}=0.2$.
4.2 freestyle 3D portrait synthesis
@@
In this section, we show the 3D portrait synthesis results of our approach. As shown in Fig. 3, our method can generate diverse style 3D portrait, and the synthesized portraits are high quality and 3D consistent, proving our ability of freestyle generation. Then, we display the outputs of different stage (i.e., EG3D, IP2P and final output) in Fig. 4, it can be seen EG3D outputs the original 3D portrait, IP2P will add the style into the portrait images (the facial details and consistency can not be guaranteed), our method will generate the final high quality and 3D-consistency results. In addition, we generate the 3D geometry results of different style in Fig. 5, which can accurately reflect the given style. Last but not least, Fig. 6 shows the smooth style mixing results of our method. We interpolate the encoded features of two different style in the $\mathbf{w}_{s}$ latent space, and use the new triplane implicit representation to synthesize 3D portraits. The results reveal the controllability and potential style editing ability of our model.